OpenStudy (anonymous):

Find: -1 + 1/3 -1/5 + 1/7 - 1/9 ....... till infinity.

OpenStudy (experimentx):

looks like expansion of arctan

OpenStudy (experimentx):

- pi/4 ?

OpenStudy (anonymous):

Thats correct. Proof?

OpenStudy (experimentx):

taylor expansion of arctan(x), x= -1

OpenStudy (anonymous):

Oh wait. its the geometric series one right?

OpenStudy (anonymous):

Oh. Riggghtttt. Now that you mentioned it, we can use the geometric series to get this too. Nicee.

OpenStudy (experimentx):

i remember mukushla ... providing a nice solution of it. kinda forget though ... try to get ... some series = arctan(x)

OpenStudy (anonymous):

Oh i got it now. Look. Using sum of infinite GP, => 1/1+x^2 = 1 -x^2 + x^4 - x^6..........

OpenStudy (experimentx):

int 1/(1+x^2) = arctan(x)

OpenStudy (anonymous):

Where -x^2 is the common ratio. Now we integrate it from 0 to 1.

OpenStudy (anonymous):

To get this sum.

OpenStudy (experimentx):

yep ... correct.

OpenStudy (anonymous):

Thanks! :D

OpenStudy (experimentx):

kinda had rev too ...

OpenStudy (experimentx):

nice prob ...